OpenStudy (anonymous):

solve the system the first equation is 3xy-4(x^2)=2 and the second equation is -5(x^2)+3y^2=7

OpenStudy (anonymous):

\[{3xy-4x^2=2,- 5x^2+3y^2=7}\]

OpenStudy (anonymous):

i need to know how to solve it for all its real solutions

OpenStudy (mertsj):

Solve the first equation for y and substitute into the second equation.

OpenStudy (anonymous):

yeah but there are 4 solutions

OpenStudy (mertsj):

Yes. And when you do that you will get an equation containing x^4 which will have 4 solutions.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

\[y=\frac{2+4x^2}{3x}\]

OpenStudy (mertsj):

Substitute that into the second equation.

OpenStudy (anonymous):

so i just solve that

OpenStudy (mertsj):

Do you know what substitute means?

OpenStudy (anonymous):

ok yes got it

OpenStudy (mertsj):

\[-5x^2+3(\frac{2+4x^2}{3x})^2=7\]

OpenStudy (mertsj):

\[-5x^2+3(\frac{4+16x^2+16x^4}{9x^2})=7\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i already have two solutions i just need to know how to get the other two

OpenStudy (mertsj):

\[-5x^2+\frac{12+48x^2+48x^4}{9x^2}=7\]

OpenStudy (mertsj):

What are the two you have?

OpenStudy (anonymous):

(-3,-3) and (2,3) and i know th other two are (-1,-2) and (1,20 i just don't know how to get them

OpenStudy (mertsj):

You end up with: \[(x^2-4)(x^2-1)=0\]

OpenStudy (anonymous):

sorry the first one is (-2,-3)

OpenStudy (mertsj):

So x = 2, -2, 1, -1

OpenStudy (mertsj):

So go back to the equation:

OpenStudy (mertsj):

\[y=\frac{2+4x^2}{3x}\]

OpenStudy (mertsj):

And replace x with each one of the 4 x values and find the corresponding y.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so r (2,3) and (-2,-3) not solutions

OpenStudy (mertsj):

(2,6) (-2,-6) (1,2) (-1,-2)

OpenStudy (anonymous):

abc123 that is true

OpenStudy (anonymous):

what is

OpenStudy (mertsj):

That (2,3) and (-2,-3) are not solutions.

OpenStudy (anonymous):

(2,3) and (-2,-3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What grade are you in bro?

OpenStudy (anonymous):

11. don't judge

OpenStudy (mertsj):

No wait. Those two are solutions. The 4 solutions are:(2,3, )(-2,-3, )(1,2), (-1,-2)

OpenStudy (mertsj):

Sorry. I made a calculation error earlier.

OpenStudy (anonymous):

ok